# Stay-Positive: A Case for Ignoring Real Image Features in Fake Image Detection

### v1

@@ 6 Limitations
Our approach enhances existing detectors by ensuring that the final layer ignores features associated with the real distribution. However, patterns linked to the fake distribution can also be spurious. For instance, Rajan et al. (2024) observed that Corvi incorrectly associates upsampled images with the real distribution. We find that Corvidirect-$\oplus$ (Ours) exhibits a similar issue. To illustrate this, we take 500 Redcaps images (512×512), upsample them, and analyze how their upsampled versions affect the probability of being classified as fake. As shown in Fig. 7, upsampling spuriously increases this probability.
Our method ensures that the detector ignores real image-specific features, but these features can still shape its notion of fakeness. For instance, the network may infer that the absence of certain real image features indicates a fake image. We illustrate this with a simple toy example in Appendix A.4. Thus, directly ignoring real features without a two-stage training process may be a promising next direction.
@@ 7 Related Work
Training-based methods create effective fake image detectors. Wang et al. (2020) demonstrated that data augmentations during training improve generalization. Odena et al. (2016) demonstrated identifiable artifacts, like checkerboard patterns, in the Fourier transforms of fake images. Building on this, Zhang et al. (2019) trained on Fourier images, leading to improvements. Chai et al. (2020) improved detection using patch-based classification. Gragnaniello et al. (2021) removed downsampling in the initial layers and applied patch-wise training for further gains. These techniques were extended to the LDM setting by Corvi et al. (2023). Additionally, works such as Chen et al. (2024) and Rajan et al. (2024) also use reconstructions of real images as part of the training data. These approaches struggle to generalize across architectures. To address this, Ojha et al. (2023) proposed using general-purpose visual encoders like CLIP (Radford et al., 2021) for fake image detection. Unlike ours, these methods explicitly rely on real features.
Contrary to prior approaches, GenDet (Zhu et al., 2023) treats fake image detection as an anomaly detection problem, similar to ours, but focuses on learning the real image distribution. Another paradigm suggests generators reconstruct fake images more easily than real ones. For reconstruction, Pasquini et al. (2023) use GAN inversion (Xia et al., 2022), DIRE (Wang et al., 2023) applies DDIM inversion (Song et al., 2020), and AEROBLADE (Ricker et al., 2024) leverages a latent diffusion autoencoder. However, these methods struggle with detecting post-processed images (Rajan et al., 2024). A contrasting approach, akin to image reconstruction, was recently proposed by Cozzolino et al. (2025), who use neural image-compression networks (Cao et al., 2020) to model real distribution likelihood, assuming fake images are less likely to be part of it. While this method models the real distribution, we argue the focus should be on detecting fake image artifacts instead.
@@ 8 Conclusion
We showed that learning patterns from the real distribution can harm fake image detection and propose a last-layer fine-tuning strategy to improve performance. By ignoring real-distribution features, the model reduces susceptibility to spurious correlations, enhancing robustness. Additionally, models trained this way excel on partially inpainted real images. We hope our study contributes to developing more robust detectors with positive societal impact.
@@ A.1 Implementation Details
We follow the training recipe used by Corvi et al. (2023). We train on 96 x 96 crops of the whole image using a batch size of 128. The data augmentations include random JPG compression and blur from the pipeline proposed by Wang et al. (2020). Following Gragnaniello et al. (2021), grayscale, cutout and random noise are also used as augmentations. Finally, in order to make the network invariant towards resizing, the random resized crop was added. For the baseline detectors as well as our re-trained variant, we report the average across two trained networks. When applying Algorithm 1, we use a batch size of 2048 to perform last-layer retraining. The rest of the training recipe does not differ from the original model.
@@ A.2.2 Results
We pass the real images through Corvidirect-$\oplus$ (Ours) and Rajandirect-$\oplus$ (Ours) and plot the probability of fakeness (output of the network) in the form of violin plots in Fig 9. We observe that the fakeness scores of our test distribution indicated by “Test Real” is extremely similar to the other distributions of real images such as GTA, Cubism etc. This shows that the test-set which we use in the main paper is representative of various different types of real image families. It is important to note that these real images are vastly different, belonging to different domains such as realistic scenes, art styles and video game rendered environments. However, the thing in common is the absence of the generator artifacts which is indicated by the low probability of fakeness in these violin plots.
A.3 Performance on GenImage
A.3.1 Experimental Setup
@@ A.3.2 Results
A.4 Spurious Correlations in the Real Distribution Could Influence Fake Features
@@ A.4.1 A Toy Example
A.5 Improved Detection of GAN-generated Images
Our work in the main paper focused on improving the detection of LDM (Rombach et al., 2022) generated images. In this section, we extend the results to the detection of images generated by GANs (Goodfellow et al., 2020). We consider a baseline ResNet-50 trained using the dataset created by Wang et al. (2020). This dataset consists of 360k real images taken from LSUN and 360k images generated by ProGAN (Karras et al., 2018). We use the same training recipe described in Appendix A.1, however, we do not employ the random resized crop data augmentation in this case. We train a ResNet-50 on this dataset. We refer to this baseline as GAN-Baseline, we re-train the last layer of this detector to ignore real features. We refer to this model as GAN-Baselinedirect-$\oplus$ (Ours).
A.5.1 Improved Robustness to WEBP Compression
@@
Experiment Details: With this experiment, we intend to measure the sensitivity of the original GAN-Baseline to WEBP compression and measure if our method can mitigate these issues. We use the StyleGAN (Karras, 2019) test set provided by Wang et al. (2020). We sample 500 real images and 500 fake images randomly. We apply different levels of WEBP compression to the fake images and compute the AP, analogous to our measurement in Section 3.1.
@@
A.5.2 General Performance
In this section, we study the generalization ability of the detector in comparison to other baseline detectors. We use the dataset consisting of CNN generated images from Wang et al. (2020). The dataset consists of ProGAN (Karras et al., 2018), StyleGAN (Karras, 2019), StyleGAN2 (Karras et al., 2020), GauGAN (Park et al., 2019), BigGAN (Brock, 2018), CycleGAN (Zhu et al., 2017), StarGAN (Choi et al., 2018) as well as Deepfakes (Rossler et al., 2019), super-resolution (SAN) (Dai et al., 2019) as well as networks which use a perceptual loss to refine images, such as CRN (Chen & Koltun, 2017) and IMLE (Li et al., 2019). It also consists of high quality GAN-generated faces in WFIR (whi, ). This test set comes with both real and fake images for each generator category. Therefore, we report both accuracy with a 0.5 threshold and the AP.
@@ A.5.2 General Performance
UFD-ProGAN 99.81 84.92 74.96 99.47 95.07 98.33 95.74 68.56 56.62 56.58 69.10 87.20
UFD-LDM 93.18 83.20 83.95 90.75 87.90 94.62 84.36 54.46 79.22 74.33 85.00 70.35
ClipDet 72.96 70.49 70.86 83.94 74.10 87.28 54.95 50.87 77.62 53.10 53.69 53.10
GAN-Baseline 100.00 99.52 93.34 95.53 96.65 93.14 99.54 55.43 54.33 99.96 99.87 100.00
GAN-Baselinedirect-$\oplus$ 100.00 99.82 98.96 94.52 95.02 93.30 99.77 68.17 75.57 86.75 86.75 99.70
@@
We report the accuracy and AP scores in Tables 5 and 6, respectively. Consistent with our findings in Table 2, we observe that the full-network training paradigm represented by GAN-Baseline and GAN-Baseline $\oplus$ (Ours) outperforms the CLIP-linear probing paradigm represented by the UFD methods and ClipDet. Our method, GAN-Baselinedirect-$\oplus$, outperforms the GAN-Baseline detector by 5.67 AP on images generated through super-resolution. Since super-resolution operates on real images, we hypothesize that the real features in these images harm the performance of the GAN-Baseline detector. However, our method is able to better detect such images since it does not rely on real features. Interestingly, our method shows lower accuracy on the CRN and IMLE test sets. Upon closer inspection, we find that the real distribution in these test sets consists of video game images from the GTA series, which differ significantly from the natural distribution. Notably, some patterns that the detector associates with fake images are also present in these video game images. To verify this, we conduct an experiment similar to the one in Section A.2. Except for the class “Test Real”, all other classes use the same images as described in Section A.2. We sample the ‘Test Real’ distribution from the real images in the StyleGAN test set. We observe that most of the real images, except for the GTA images, exhibit similar fakeness scores. This shows that our model still does a good job in generalizing to different domains of real images, however, the performance on some unrelated domains can drop.

### v2

@@ 6 Limitations
Our approach enhances existing detectors by ensuring that the final layer ignores features associated with the real distribution. However, patterns linked to the fake distribution can also be spurious. For instance, Rajan et al. (2025) observed that Corvi incorrectly associates upsampled images with the fake distribution. We find that Corvidirect-$\oplus$ (Ours) exhibits a similar issue. To illustrate this, we take 500 Redcaps images (512×512), upsample them, and analyze how their upsampled versions affect the logit score. As shown in Fig. 7, we can see that upsampled images are more likely to be classified as fake. Despite the improvements we demonstrate, we emphasize that greater care must be taken when curating the set of real and fake images to avoid such behavior.
Our method encourages the detector to ignore features that are specific to real images. However, after the first stage of training, these features can still implicitly influence the detector’s understanding of what constitutes a fake. For example, the model may learn that the absence of certain real-specific cues is indicative of a fake image, effectively using real image features in a negative sense. We observe such behavior in our own detector. To illustrate this effect, we include a simple toy example in Appendix A.6. These limitations suggest that better generalization could be achieved by extending this approach to train the entire network, as opposed to just the final layer.
@@ 7 Related Work
Training-based methods create effective fake image detectors. Wang et al. (2020) demonstrated that data augmentations during training improve generalization. Odena et al. (2016) demonstrated identifiable artifacts, like checkerboard patterns, in the Fourier transforms of fake images. Building on this, Zhang et al. (2019) trained on Fourier images, leading to improvements. Chai et al. (2020) improved detection using patch-based classification. Gragnaniello et al. (2021) removed downsampling in the initial layers and applied patch-wise training for further gains. These techniques were extended to the LDM setting by Corvi et al. (2023). Additionally, works such as Chen et al. (2024a) and Rajan et al. (2025) also use reconstructions of real images as part of the training data. These approaches struggle to generalize across architectures. To address this, Ojha et al. (2023) proposed using general-purpose visual encoders like CLIP (Radford et al., 2021) for fake image detection. Unlike ours, these methods explicitly rely on real features.
Contrary to prior approaches, GenDet (Zhu et al., 2023) treats fake image detection as an anomaly detection problem, similar to ours, but focuses on learning the real image distribution. Another paradigm suggests generators reconstruct fake images more easily than real ones. For reconstruction, Pasquini et al. (2023) use GAN inversion (Xia et al., 2022), DIRE (Wang et al., 2023), ZeroFake (Sha et al., 2024) apply DDIM inversion (Song et al., 2021), and AEROBLADE (Ricker et al., 2024) leverages a latent diffusion autoencoder. However, these methods at times struggle with detecting post-processed images (Rajan et al., 2025). A contrasting approach, akin to image reconstruction, was recently proposed by Cozzolino et al. (2025), who use neural image-compression networks (Cao et al., 2020) to model real distribution likelihood, assuming fake images are less likely to be part of it. While this method models the real distribution, we argue the focus should be on detecting fake image artifacts instead.
In the literature on mitigating spurious correlations, two-stage training strategies have been previously proposed. Liu et al. (2021) select samples with high training loss in the first stage and re-train the network on these samples in the second stage. Kirichenko et al. (2023) employ a last-layer re-training in a strategy. However, while these approaches rely on data curation in the second stage to force the network to focus on relevant features, our approach instead constrains the network architecture itself, encouraging it to focus on patterns that are present in fake images.
@@ 8 Conclusion
We showed that learning patterns from the real distribution can harm fake image detection and propose a last-layer fine-tuning strategy to improve performance. By ignoring real-distribution features, the model reduces susceptibility to spurious correlations, enhancing robustness. Additionally, models trained this way excel on partially inpainted real images. While this work focuses on detecting AI-generated images, we believe that our core idea — that specific patterns should not be associated with the real distribution could be applicable to other forms of media forensics, such as audio and video. We leave these directions to future research. We hope our study contributes to developing more robust detectors with positive societal impact.
@@ A.1 Implementation Details
We follow the training recipe used by Corvi et al. (2023). We train on 96 x 96 crops of the whole image using a batch size of 128. The data augmentations include random JPG compression and blur from the pipeline proposed by Wang et al. (2020). Following Gragnaniello et al. (2021), grayscale, cutout and random noise are also used as augmentations. Finally, in order to make the network invariant towards resizing, the random resized crop was added. For the baseline detectors as well as our re-trained variant, we report the average across two trained networks. When applying Algorithm 1, we use a batch size of 1024 to perform last-layer retraining. The rest of the training recipe does not differ from the original model. The second stage training on average converges in about 15 epochs, which takes an additional 4 hours.
@@ A.2.2 Results
We pass the real images through Corvidirect-$\oplus$ (Ours) and Rajandirect-$\oplus$ (Ours) and plot the logit score (output of the network) in the form of violin plots in Fig 9. We observe that the fakeness scores of our test distribution indicated by “Test Real” is extremely similar to the other distributions of real images such as GTA, Cubism etc. This shows that the test-set which we use in the main paper is representative of various different types of real image families. It is important to note that these real images are vastly different, belonging to different domains such as realistic scenes, art styles and video game rendered environments. However, the thing in common is the absence of the generator artifacts which is indicated by the low logit score in these violin plots.
A.3 Robustness Analysis
In the main paper, we show that our method shows improved robustness to WEBP Compression and downsizing. In this section, we test the robustness of our detector to JPEG Compression, additive gaussian noise and low-pass filtering. For JPEG and additive noise, we follow the same experimental setting adopted by Rajan et al. (2025). Instead of AP, we report the logit scores. This way, we study the behaviour of the detector on both post-processed real and fake images,
(a) JPEG Compression
(b) Additive Gaussian Noise
(c) Low-Pass Filtering
Figure 10: Robustness to Common Post-Processing Artifacts. We observe similar trends across various corruptions (JPEG compression, additive noise, and low-pass filtering) between the original detectors and our improved versions. Note that these perturbations were part of the training data augmentations. More importantly, the real image distribution exhibits very uniform logit scores, since our approach relies on the absence of generator artifacts to identify these images.
We report the results from robustness tests in Fig 10. We can observe that our variants perform similar to the original models, which is not surprising since all of these perturbations were seen during training. Additionally, our real distribution shows very little variance unlike the distribution of the original baselines since our method does not rely on specific features to call an image real. Additionally, we also test the robustness to resizing using the SynthBuster dataset (akin to Fig 8 from Cozzolino et al. (2024)) and we report our results in Fig 11.
Figure 11: Robustness to Resizing (SynthBuster) We can observe that our detector, Corvidirect-$\oplus$ shows improved robustness to downsizing compared to the original Corvi
A.4 Performance on GenImage
A.4.1 Experimental Setup
@@ A.4.2 Results
A.5 Performance on the UFD Benchmark
A.5.1 Experimental Setup
In this section, we evaluate Corvi $\oplus$ (Ours) and Rajan $\oplus$ (Ours) on pixel-space diffusion models such as ADM (Dhariwal & Nichol, 2021) and GLIDE (Nichol et al., 2022), as well as on autoregressive models such as DALL-E (Ramesh et al., 2021) and Stable Diffusion (Rombach et al., 2022). Since our focus is on comprehensive detection of images from a known generator family, we do not consider GAN-generated images in this evaluation.
Table 6: UFD Benchmark AP Scores across various diffusion models and detectors. Detectors trained with our improved algorithm (Corvidirect-$\oplus$, Rajandirect-$\oplus$) consistently outperform their base versions and other models across all UFD benchmark settings, demonstrating superior generalization and robustness.
GLIDE ADM DALLE LDM AVG
Method 100-27 100-10 50-27 100 200 200-cfg
Corvi 79.28 86.52 83.95 44.51 97.15 99.90 100.00 99.99 86.41
Corvidirect-$\oplus$ (Ours) 88.61 91.17 89.24 74.52 99.50 100.00 100.00 99.99 92.88
Rajan 71.51 77.57 77.54 44.63 89.19 99.90 100.00 99.99 82.54
Rajandirect-$\oplus$ (Ours) 80.43 83.50 83.36 62.32 98.63 100.00 100.00 99.99 88.53
A.5.2 Results
We report the results in Table 6. Both Corvidirect-$\oplus$ and Rajandirect-$\oplus$ demonstrate substantial improvements over their original counterparts in terms of AP on images generated by GLIDE, ADM, and DALL-E. These results indicate that, even when generalizing to entirely unseen settings, disregarding real-image features remains an effective strategy.
A.6 Spurious Correlations in the Real Distribution Could Influence Fake Features
@@ A.6.1 A Toy Example
A.7 Improved Detection of GAN-generated Images
Our work in the main paper focused on improving the detection of LDM (Rombach et al., 2022) generated images. In this section, we extend the results to the detection of images generated by GANs (Goodfellow et al., 2014). We consider a baseline ResNet-50 trained using the dataset created by Wang et al. (2020). This dataset consists of 360k real images taken from LSUN and 360k images generated by ProGAN (Karras et al., 2018). We use the same training recipe described in Appendix A.1, however, we do not employ the random resized crop data augmentation in this case. We train a ResNet-50 on this dataset. We refer to this baseline as GAN-Baseline, we re-train the last layer of this detector to ignore real features. We refer to this model as GAN-Baselinedirect-$\oplus$ (Ours).
A.7.1 Improved Robustness to WEBP Compression
@@
Experiment Details: With this experiment, we intend to measure the sensitivity of the original GAN-Baseline to WEBP compression and measure if our method can mitigate these issues. We use the StyleGAN (Karras et al., 2019) test set provided by Wang et al. (2020). We sample 500 real images and 500 fake images randomly. We apply different levels of WEBP compression to the fake images and compute the AP, analogous to our measurement in Section 3.1.
@@
A.7.2 General Performance
In this section, we study the generalization ability of the detector in comparison to other baseline detectors. We use the dataset consisting of CNN generated images from Wang et al. (2020). The dataset consists of ProGAN (Karras et al., 2018), StyleGAN (Karras et al., 2019), StyleGAN2 (Karras et al., 2020), GauGAN (Park et al., 2019), BigGAN (Brock et al., 2019), CycleGAN (Zhu et al., 2017), StarGAN (Choi et al., 2018) as well as Deepfakes (Rossler et al., 2019), super-resolution (SAN) (Dai et al., 2019) as well as networks which use a perceptual loss to refine images, such as CRN (Chen & Koltun, 2017) and IMLE (Li et al., 2019). It also consists of high quality GAN-generated faces in WFIR (whi, ). This test set comes with both real and fake images for each generator category. Therefore, we report both accuracy with a 0.5 threshold and the AP.
@@ A.7.2 General Performance
Method ProGAN StyleGAN StyleGAN2 GauGAN BigGAN CycleGAN StarGAN Deepfake SAN CRN IMLE WFCIR AVG
UFD-ProGAN 99.81 84.92 74.96 99.47 95.07 98.33 95.74 68.56 56.62 56.58 69.10 87.20 82.20
UFD-LDM 93.18 83.20 83.95 90.75 87.90 94.62 84.36 54.46 79.22 74.33 85.00 70.35 81.78
ClipDet 72.96 70.49 70.86 83.94 74.10 87.28 54.95 50.87 77.62 53.10 53.69 53.10 66.91
GAN-Baseline 100.00 99.52 93.34 95.53 96.65 93.14 99.54 55.43 54.33 99.96 99.87 100.00 90.61
GAN-Baselinedirect-$\oplus$ 100.00 99.82 98.96 94.52 95.02 93.30 99.77 68.17 75.57 86.75 86.75 99.70 91.53
@@
We report the accuracy and AP scores in Tables 7 and 8, respectively. Consistent with our findings in Table 2, we observe that the full-network training paradigm represented by GAN-Baseline and GAN-Baseline $\oplus$ (Ours) outperforms the CLIP-linear probing paradigm represented by the UFD methods and ClipDet. Our method, GAN-Baselinedirect-$\oplus$, outperforms the GAN-Baseline detector by 5.67 AP on images generated through super-resolution. Since super-resolution operates on real images, we hypothesize that the real features in these images harm the performance of the GAN-Baseline detector. However, our method is able to better detect such images since it does not rely on real features. Interestingly, our method shows lower accuracy on the CRN and IMLE test sets. Upon closer inspection, we find that the real distribution in these test sets consists of video game images from the GTA series, which differ significantly from the natural distribution. Notably, some patterns that the detector associates with fake images are also present in these video game images. To verify this, we conduct an experiment similar to the one in Section A.2. Except for the class “Test Real”, all other classes use the same images as described in Section A.2. We sample the ‘Test Real’ distribution from the real images in the StyleGAN test set. We observe that most of the real images, except for the GTA images, exhibit similar logit scores. This shows that our model still does a good job in generalizing to different domains of real images, however, the performance on some unrelated domains can drop.